One cow is tied per tree and three cow's are left and if two cow's are tied per tree then four trees are left. How many trees and cow's are there in total.
Answer by edjones(8007)

let c=number of cows, t= number of trees
there are 3 more cows than trees: c=t+3
when we effectively cut the number of cows in half by tying 2 to a tree we have 4 more trees than cows : c/2+4=t
.
c=t+3
c/2+4=t
.
multiply 2 times the 2nd equation: c+8=2t
replace c with t+3 in the second equation: t+3+8=2t
subtract t from both sides: t=11 Ans.
therefore c=14 Ans
Ed